the hypotenuse of a right triangle is 7 inches longer than the base and 14 inches longer than te height the perimeter of the triangle is 84 inches find its area
Answer by lwsshak3(11628)

the hypotenuse of a right triangle is 7 inches longer than the base and 14 inches longer than te height the perimeter of the triangle is 84 inches find its area
***
let x=length of hypotenuse
base,b=x-7
height=x-14
x+x-7+x-14=84
3x-21=84
3x=105
x=35
area=1/2*base*height=1/2(x-7)(x-14)=1/2(28)(21)=294 sq inches